INFRARED SPECTROSCOPY OF PROTONATED MIXED BENZENE-WATER CLUSTERS

\underline{T. CHENG}, B. BANDYOPADHYAY and M. A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602.

Mixed clusters of protonated benzene and water are created via arc discharge in a molecular beam cluster source. Infrared spectroscopy (1000 cm^{−1} to 4500 cm^{−1}) of these mixed clusters H^{+}(H_{2}O)_{x}(Bz)_{y} (x=1-4, y=1-4) tagged with argon is employed to investigate the structures of these clusters, particularly with regards to the location of the proton. Studies as a function of cluster size investigate solvation effects within the mixed clusters.